OpenStudy (anonymous):

is the following statement false 2+(12)>2(12)

OpenStudy (anonymous):

yes, because 14>24 is not correct. if the sign was the other way it would be true,but given its >, then no

OpenStudy (anonymous):

@GotMath7- Thank you, The reason I asked because this is an example question (not the real problem I need to solve) I just started a math class for the 1st time since 1998 and I admit I'm a bit more than intimated. I guess my Real question is in this instance ( the 1st ex) 2 + (12) = 14 I'm confused as to why? I thought we multiplied when we saw a Parenthesis ( )

OpenStudy (anonymous):

you usually do,but since there is nothing beside 12, i assume you would believe it to be multiplied by 1, therefore making it 2+(12)1=14, which would be 2+12=14, 14=14

OpenStudy (anonymous):

great! Thank you so much! @GotMath7